# Daria Sushnikova    George Turkiyyah    Edmond Chow    David Keyes

### v1

@@ Abstract
This paper presents a new fast iterative solver for large systems involving kernel matrices. Advantageous aspects of $\mathcal{H}^{2}$ matrix approximations and the multigrid method are hybridized to create the $\mathcal{H}^{2}$-MG algorithm. This combination provides the time and memory efficiency of $\mathcal{H}^{2}$ operator representation along with the rapid convergence of a multilevel method. We describe how $\mathcal{H}^{2}$-MG works, show its linear complexity, and demonstrate its effectiveness on two standard kernels. The current zoo of $\mathcal{H}^{2}$ solvers, which includes a wide variety of iterative and direct solvers, so far lacks a method that exploits multiple levels of resolution, commonly referred to in the iterative methods literature as “multigrid” from its origins in a hierarchy of grids used to discretize differential equations. This makes $\mathcal{H}^{2}$-MG a valuable addition to the collection of $\mathcal{H}^{2}$ solvers. The algorithm has potential for advancing various fields that require the solution of large, dense, symmetric positive definite matrices.
@@
$\mathcal{H}^{2}$matrix, multigrid methods, kernel matrices, rank-structured matrices, iterative solvers, linear complexity
@@ 1 Introduction
Rank-structured methods, and particularly $\mathcal{H}^{2}$  [20, 5] matrices, typically provide a time- and memory-efficient matrix-by-vector product, which naturally leads to solving systems using iterative solvers. However, iterative solvers have their disadvantages, as their efficiency depends on the number of iterations and thus on the matrix conditioning. Direct solvers have their own challenges, being extremely complex for rank-structured matrices and involving a large constant overhead. To fill this gap, we looked at the multigrid [12, 6, 17] method, which is typically used for sparse matrices, and adapted it for $\mathcal{H}^{2}$ matrices. Multigrid methods exhibit excellent convergence properties and can significantly benefit from fast $\mathcal{H}^{2}$ matrix-vector products. In this paper, we introduce a new algorithm, $\mathcal{H}^{2}$- multigrid ($\mathcal{H}^{2}$-MG ), which leverages the hierarchical structure of $\mathcal{H}^{2}$ matrices to create a multigrid method that operates across different levels of the $\mathcal{H}^{2}$ matrix.
@@
  • Developing the $\mathcal{H}^{2}$-MG algorithm, a hybrid solver combining multilevel resolution with hierarchical matrix representation.
  • Theoretically demonstrating the linear complexity of the algorithm in both time and memory.
  • Validating its effectiveness on two standard kernels and comparing its performance with existing approaches.
@@ 2 Related work
A general $N$ by $N$ matrix requires $\mathcal{O}(N^{2})$ operations to compute a matrix-vector product and $\mathcal{O}(N^{2})$ memory for storage. However, significant progress has been made to reduce this computational cost with controllable loss of accuracy. Block low-rank methods, such as the mosaic skeleton [35, 36] and the $\mathcal{H}$-matrix [21, 18, 25] techniques, reduce the number of operations for matrix-vector products and storage to $\mathcal{O}(N\log N)$ by exploiting low-rank approximations of certain blocks of the matrix. Building upon these advancements, nested-basis methods like HSS matrices [39, 8, 13], HODLR matrices [1], $\mathcal{H}^{2}$ matrices [20, 5, 27], and the Fast Multipole Method (FMM) [15, 40, 14] were proposed. These methods enable matrix-vector multiplication with high accuracy in $\mathcal{O}(N)$ operations for many matrices arising in physically causal models where interactions decay smoothly with distance. HSS and HODLR methods are particularly efficient for 1D problems, while $\mathcal{H}^{2}$ and FMM extend their efficiency to 2D and 3D problems thanks to their strong admissibility property.
@@
This accelerated matrix-vector multiplication becomes the basis for solving linear systems using iterative techniques such as GMRES [30], CG [22, 33], BiCGstab [37], and any other iterative solvers. Iterative methods, while versatile, have a drawback: their convergence rate depends on the conditioning or eigenvalue clustering of the matrix. In contrast, fast direct solvers guarantee a predefined number of operations for solving the system. Although the complexity for general dense matrices remains $\mathcal{O}(N^{3})$, leveraging the $\mathcal{H}$-matrix format has led to breakthroughs in direct solver efficiency. Researchers in [19] introduced an $\mathcal{O}(N\log N)$ direct solver algorithm, albeit with a substantial constant. Subsequent works [2, 28, 34] utilizing $\mathcal{H}^{2}$ (FMM) format have achieved $\mathcal{O}(N)$ direct solver algorithms with more favorable constants.
@@
In this paper, we expand the zoo of $\mathcal{H}^{2}$ solvers by introducing a novel iterative algorithm – the $\mathcal{H}^{2}$-MG solver. This solver is rooted in the standard multigrid method but tailored for $\mathcal{H}^{2}$ structures.
@@
Multigrid (MG) methods, as well as rank structured methods, play a crucial role in solving large linear systems, particularly systems with sparse matrices arising from discretized partial differential equations. Multigrid has its origins in [12] and was further developed in [6]. Multigrid has become a cornerstone of numerical techniques for solving large sparse linear systems. Its strategy lies in efficiently reducing errors at multiple scales by leveraging a hierarchy of coarser grids. Over time, a variety of fruitful generalizations have emerged, such as algebraic multigrid (AMG) [7] approaches that extend its applicability to unstructured grids and irregular geometries [11, 9, 16] and fast multipole preconditioners for sparse matrices [23].
@@
Our approach of applying multigrid to the system with $\mathcal{H}^{2}$ matrix diversifies the landscape of $\mathcal{H}^{2}$ solvers, offering a promising alternative that is relatively easier to implement and parallelize efficiently compared to direct solvers. Demonstration of parallel scaling, however, is beyond the scope of this initial description.
@@ 3.1 $\mathcal{H}^{2}$ matrix
In this section, we briefly review the fundamental concept of an $\mathcal{H}^{2}$ matrix, a hierarchical block low-rank matrix structure with a nested basis property. Matrices well approximated in $\mathcal{H}^{2}$ form typically come from the discretization of boundary integral equations and several other problems with approximately separable kernels. The hierarchical nature of $\mathcal{H}^{2}$ matrix is the basis and main inspiration of the $\mathcal{H}^{2}$-MG algorithm. For more comprehensive and formal $\mathcal{H}^{2}$ definitions, see [15, 20, 5].
@@
and is called a “close” block, or has a low rank, is called “far”, and possess the following property:
@@
where $\widehat{D}_{ij}\in\mathbb{R}^{B_{i}\times B_{j}}$, $\widehat{U}_{i}\in\mathbb{R}^{B_{i}\times r_{i}}$, $\widehat{S}_{ij}\in\mathbb{R}^{r_{i}\times r_{j}}$, $\widehat{V}_{j}\in\mathbb{R}^{r_{j}\times B_{j}}$. Note that all low-rank blocks in a row $i$ have the same left factor $\widehat{U}_{i}$, and all the low-rank blocks in a column $j$ have the same right factor $\widehat{V}_{j}$. This is one of the defining features of the $\mathcal{H}^{2}$ matrix. We denote $r_{i}$ as the rank of $i$-th block row, excluding full-rank blocks, and $r_{j}$ is the rank of $j$-th block column, excluding full-rank blocks.
@@
To write equation (1) in a matrix form, we define rectangular diagonal matrices $U_{1}\in\mathbb{R}^{N\times N_{2}}$, $V_{1}\in\mathbb{R}^{N_{2}\times N}$, where ${N_{2}=\sum_{i=1}^{M}r_{i}}$.
$U_{1}=\begin{bmatrix}\widehat{U}_{1}&&\\ &\ddots&\\ &&\widehat{U}_{M}\end{bmatrix},\quad V_{1}=\begin{bmatrix}\widehat{V}_{11}&&\\ &\ddots&\\ &&\widehat{V}_{MM}\end{bmatrix},$
Also, define a matrix $S_{1}\in\mathbb{R}^{N_{2}\times N_{2}}$:
$[S]_{ij}=\begin{cases}0,&\quad\text{if }A_{ij}\text{ is a close block}\\ \widehat{S}_{ij},&\quad\text{if }A_{ij}\text{ is a far block}\end{cases}.$
These definitions allow us to rewrite equation (1) in a matrix form:
$F=U_{1}S_{1}V_{1}$
The equation (2) then evolves to:
@@
This decomposition is illustrated in Figure 1.
@@
The figure illustrates a special case; the matrix $D$ may have more complex block structure. In general, it could be any block-sparse matrix.
@@
where $D_{2}\in\mathbb{R}^{N_{2}\times N_{2}}$, $U_{2}\in\mathbb{R}^{N_{2}\times N_{3}}$, $V_{2}\in\mathbb{R}^{N_{3}\times N_{2}}$, $S_{2}\in\mathbb{R}^{N_{3}\times N_{3}}$. The size of $N_{3}$ is $\sum_{i=1}^{\frac{M}{p}}r_{i}$, where $r_{i}$ are the ranks of the far blocks of the matrix $S_{1}$. An illustration of the factorization (3) is in Figure 2.
@@
Denote $x_{i}=V_{i}x_{i-1}$, $x_{i}\in\mathbb{R}^{N_{i}}$, $i=1\dots l$, assuming $x=x_{0}$, to obtain:
@@
Denote $y_{i-1}=D_{i}x_{i-1}+U_{i}y_{i}$, $i=1,\dots,l$, $y_{l}=S_{l}x_{l}$, and note that that $D_{1}=D$, $y_{0}=y$, to obtain:
@@
For the $\mathcal{H}^{2}$ matrix $A\in\mathbb{R}^{N\times N}$, according to [5], the storage requirement scales as $\mathcal{O}(N)$. $\mathcal{H}^{2}$ construction and matrix-by-vector multiplication complexity is also $\mathcal{O}(N)$.
@@ 3.2 Multigrid method
A multigrid method [6, 17] is a powerful numerical technique used for solving partial differential equations (PDEs) and linear systems of equations with sparse matrices. While a solver in its own right, it is often used to precondition other iterative algorithms to cause them to converge rapidly.
@@ 3.3 $\mathcal{H}^{2}$-MG
In this section, we implement a multigrid algorithm for a system with $\mathcal{H}^{2}$ matrix. Consider the system
@@
In our numerical experiments, which are on symmetric positive definite matrices, we use CG [33, 22] solver as a smoother because of its fast and easily parallelized application. Thanks to the $\mathcal{H}^{2}$ structure of matrix $A$, the application of one iteration of CG is linear in time and memory cost. After smoothing, we receive an approximation of the error:
@@
Then, we compute the residual:
$\hat{r}_{1}=r_{1}-A\tilde{e}_{1}.$
Then we subtract the term $A\tilde{e}_{1}$ from both parts of equation (6),
@@
Denoting $\hat{e}_{1}=e_{1}-\tilde{e}_{1}$ to obtain the system
@@
Now is the time to build the restriction and prolongation operators. If we denote the matrix $A$ explicitly in its $\mathcal{H}^{2}$ format, we obtain an expanded version of the equation (7):
@@
Then, reduce all identity matrices:
@@
We denote the restricted operator by $A_{2}$:
(8) $A_{2}=U^{\top}_{1}DV^{\top}_{1}+D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2}).$
@@
Note that $A_{2}\in\mathbb{R}^{N_{2}\times N_{2}}$ has $\mathcal{H}^{2}$ structure, just like matrix $A$. The matrices $A$ and $A_{2}$ have exactly the same structure, except that matrix $A$ stores the full matrix $D$, while $A_{2}$ stores a reduced part of it, $U_{1}^{\top}DV^{\top}$. Thus, matrix $A_{2}$ stores less information than $A$. Since $A$ scales linearly as an $\mathcal{H}^{2}$ matrix, $A_{2}$ also scales linearly. See the detailed proof in Section 3.4.
Denote the restricted error and residual vectors:
@@
$A_{2}e_{2}=r_{2}$
Then, according to the multigrid algorithm, we apply smoother to the restricted system and obtain the approximation of the error:
@@
where $A_{l}$ is a small dense matrix since $l$ is the top level of $\mathcal{H}^{2}$ hierarchy. We solve the system directly with dir_sol. In our computations, we use the Cholesky factorization:
@@
Then, we apply the smoothing operator starting with initial guess $\tilde{e}_{l-1}$:
$e_{l-1}=\textbf{Iter}(A_{l-1},r_{l-1},\tilde{e}{l-1}).$
@@
$\tilde{e}_{1}=\textbf{Iter}(A,r1,0)$$e_{1}=\textbf{Iter}(A,r_{1},\tilde{e}_{1})$$r_{2}=U_{1}^{\top}(r_{1}-A\tilde{e}_{1})$$r_{3}=U_{2}^{\top}(r_{2}-A_{2}\tilde{e}_{2})$$\tilde{e}_{1}=\tilde{e}_{1}+V_{1}^{\top}e_{2}$$\tilde{e}_{2}=\tilde{e}_{2}+V_{2}^{\top}e_{3}$$\tilde{e}_{2}=\textbf{Iter}(A_{2},r_{2},0)$$e_{2}=\textbf{Iter}(A_{2},r_{2},\tilde{e}_{2})$$e_{3}=\textbf{dir\_sol}(A_{3},r_{3})$
@@
The visualization of $\mathcal{H}^{2}$-MG V-cycle, presented in Figure 4, is flipped upside-down to emphasize the analogy of $\mathcal{H}^{2}$-MG method with the $\mathcal{H}^{2}$ structure. Compare $\mathcal{H}^{2}$-MG V-cycle in Figure 4 and the $\mathcal{H}^{2}$ matrix by vector product in Figure 3.

### v2

@@ Abstract
This paper presents a new fast iterative solver for large systems involving kernel matrices. Advantageous aspects of $\mathcal{H}^{2}$ matrix approximations and the multigrid method are hybridized to create the $\mathcal{H}^{2}$-MG algorithm. This combination provides the time and memory efficiency of $\mathcal{H}^{2}$ operator representation along with the rapid convergence of a multilevel method. We describe how $\mathcal{H}^{2}$-MG works, show its linear complexity, and demonstrate its effectiveness on two standard kernels and on a single-layer potential boundary element discretization with complex geometry. The current zoo of $\mathcal{H}^{2}$ solvers, which includes a wide variety of iterative and direct solvers, so far lacks a method that exploits multiple levels of resolution, commonly referred to in the iterative methods literature as “multigrid” from its origins in a hierarchy of grids used to discretize differential equations. This makes $\mathcal{H}^{2}$-MG a valuable addition to the collection of $\mathcal{H}^{2}$ solvers. The algorithm has potential for advancing various fields that require the solution of large, dense, symmetric positive definite matrices.
@@
$\mathcal{H}^{2}$-matrix, multigrid methods, kernel matrices, rank-structured matrices, iterative solvers, linear complexity
@@ 1 Introduction
Rank-structured methods, and particularly $\mathcal{H}^{2}$ matrices [23, 6], typically provide a time- and memory-efficient matrix-vector product [7], which naturally leads to solving systems of equations using iterative solvers. However, iterative solvers have their disadvantages, as their efficiency depends on the number of iterations and thus on the matrix conditioning. Direct solvers have their own challenges, being extremely complex for rank-structured matrices and involving a large constant overhead. To fill this gap, we consider the multigrid method [15, 9, 20], which is typically used for sparse matrices, and adapt it to $\mathcal{H}^{2}$ matrices. Multigrid methods exhibit excellent convergence properties and can significantly benefit from fast $\mathcal{H}^{2}$ matrix-vector products. In this paper, we introduce a new algorithm, $\mathcal{H}^{2}$-multigrid ($\mathcal{H}^{2}$-MG), which leverages the hierarchical structure of $\mathcal{H}^{2}$ matrices to create a multigrid method that operates across different levels of the $\mathcal{H}^{2}$ matrix.
@@
  • Developing the $\mathcal{H}^{2}$-MG algorithm, a hybrid solver combining multilevel resolution with a hierarchical matrix representation.
  • Demonstrating the linear complexity of $\mathcal{H}^{2}$-MG in both time and memory.
  • Validating the effectiveness of $\mathcal{H}^{2}$-MG on problems from two standard kernel functions and a boundary element method, and comparing its performance with existing approaches.
@@ 2 Related work
A general $N$ by $N$ matrix requires $\mathcal{O}(N^{2})$ operations to compute a matrix-vector product and $\mathcal{O}(N^{2})$ memory for storage. However, significant progress has been made to reduce this computational cost with controllable loss of accuracy. Block low-rank matrix representations, such as the mosaic skeleton [38, 39], $\mathcal{H}$-matrices [24, 21, 28], and HODLR matrices [1], reduce the number of operations for matrix-vector products and storage to $\mathcal{O}(N\log N)$ by exploiting low-rank approximations of certain blocks of the matrix. Further, nested-basis representations like HSS matrices [42, 11, 16], $\mathcal{H}^{2}$ matrices [23, 6, 30], and the Fast Multipole Method (FMM) [18, 44, 17] were proposed. These methods enable matrix-vector multiplication with high accuracy in $\mathcal{O}(N)$ operations for many matrices arising in physically causal models where interactions decay smoothly with distance [7, 8]. HSS and HODLR methods are particularly efficient for 1D problems, while $\mathcal{H}^{2}$ and FMM extend their efficiency to 2D and 3D problems thanks to their strong admissibility property.
@@
This accelerated matrix-vector multiplication becomes the basis for solving linear systems using iterative techniques such as GMRES [33], CG [25, 36], BiCGstab [40], and other iterative solvers. Iterative methods, while versatile, have a drawback: their convergence rate depends on the conditioning or eigenvalue clustering of the matrix. Thus, preconditioning for iterative methods, where the matrix is in $\mathcal{H}^{2}$matrix format, is often a necessity [43, 45]. In contrast, fast direct solvers guarantee a predefined number of operations for solving the system. Although the complexity for general dense matrices remains $\mathcal{O}(N^{3})$, leveraging the hierarchical matrix format has led to breakthroughs in direct solver efficiency. Researchers in [22] introduced an $\mathcal{O}(N\log N)$ direct solver algorithm, albeit with a substantial constant. Subsequent works [2, 31, 37] utilizing $\mathcal{H}^{2}$ (FMM) format have achieved $\mathcal{O}(N)$ direct solver algorithms with more favorable constants.
@@
In this paper, we expand the zoo of $\mathcal{H}^{2}$ solvers by introducing a novel iterative algorithm—the $\mathcal{H}^{2}$-MG solver. This solver is rooted in the standard multigrid method but tailored for $\mathcal{H}^{2}$ structures.
@@
Multigrid (MG) methods, as well as rank structured methods, play a crucial role in solving large linear systems, particularly systems with sparse matrices arising from discretized partial differential equations. Multigrid has its origins in [15] and was further developed in [9]. Multigrid has become a cornerstone of numerical techniques for solving large sparse linear systems. Its strategy lies in efficiently reducing errors at multiple scales by leveraging a hierarchy of coarser grids. Over time, a variety of fruitful generalizations have emerged, such as algebraic multigrid (AMG) [10] approaches that extend its applicability to unstructured grids and irregular geometries [14, 12, 19] and fast multipole preconditioners for sparse matrices [26]. Multigrid methods have also been turned into $\mathcal{H}^{2}$ matrices in order to provide fast methods for evaluating integral operators [5].
@@
Our approach of applying multigrid to the $\mathcal{H}^{2}$ matrix diversifies the landscape of $\mathcal{H}^{2}$ solvers, offering a promising alternative that is easier to implement and parallelize efficiently compared to direct solvers. Demonstration of parallel scaling, however, is beyond the scope of this initial description.
@@ 3.1 $\mathcal{H}^{2}$ matrix
In this section, we briefly review the fundamental concept of an $\mathcal{H}^{2}$ matrix, a hierarchical block low-rank matrix structure with a nested basis property. Matrices well approximated in $\mathcal{H}^{2}$ form typically come from the discretization of boundary integral equations and several other problems with approximately separable kernels. The hierarchical nature of $\mathcal{H}^{2}$ matrices is the main inspiration for the $\mathcal{H}^{2}$-MG algorithm. For comprehensive and formal $\mathcal{H}^{2}$ definitions, see [18, 23, 6].
@@
and is called a “close” block, or has a low rank, is called “far”, and possesses the following property:
@@
with $\widehat{U}_{i}\in\mathbb{R}^{B_{i}\times r_{i}}$, $\widehat{S}_{ij}\in\mathbb{R}^{r_{i}\times r_{j}}$, $\widehat{V}_{j}\in\mathbb{R}^{r_{j}\times B_{j}}$. Note that all low-rank blocks in a row $i$ have the same left factor $\widehat{U}_{i}$, and all the low-rank blocks in a column $j$ have the same right factor $\widehat{V}_{j}$. This is one of the defining features of the $\mathcal{H}^{2}$ matrix. We denote $r_{i}$ as the rank of $i$-th block row, excluding full-rank blocks, and $r_{j}$ is the rank of $j$-th block column, excluding full-rank blocks.
@@
To write equation (1) in matrix form, we define rectangular diagonal matrices $U_{1}\in\mathbb{R}^{N\times N_{2}}$, $V_{1}\in\mathbb{R}^{N_{2}\times N}$, where ${N_{2}=\sum_{i=1}^{M}r_{i}}$:
$U_{1}=\begin{bmatrix}\widehat{U}_{1}&&\\ &\ddots&\\ &&\widehat{U}_{M}\end{bmatrix},\quad V_{1}=\begin{bmatrix}\widehat{V}_{1}&&\\ &\ddots&\\ &&\widehat{V}_{M}\end{bmatrix},$
We also define a matrix $S_{1}\in\mathbb{R}^{N_{2}\times N_{2}}$ as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}[S_{1}]_{ij}}=\begin{cases}0,&\quad\text{if }A_{ij}\text{ is a close block}\\ \widehat{S}_{ij},&\quad\text{if }A_{ij}\text{ is a far block}\end{cases}.$
These definitions allow us to rewrite equation (1) in matrix form as:
$F=U_{1}S_{1}V_{1},$
which allows us to express equation (2) as:
@@
This decomposition is illustrated in Figure 1. The figure illustrates a special case; the matrix $D$ may have more complex block structure. In general, it could be any block-sparse matrix.
@@
with $D_{2}\in\mathbb{R}^{N_{2}\times N_{2}}$, $U_{2}\in\mathbb{R}^{N_{2}\times N_{3}}$, $V_{2}\in\mathbb{R}^{N_{3}\times N_{2}}$, $S_{2}\in\mathbb{R}^{N_{3}\times N_{3}}$. $N_{3}$ is $\sum_{i=1}^{\frac{M}{p}}r_{i}$, where $r_{i}$ are the ranks of the far blocks of the matrix $S_{1}$. An illustration of the factorization (3) is shown in Figure 2.
@@
Letting $x_{i}=V_{i}x_{i-1}$, $x_{i}\in\mathbb{R}^{N_{i}}$, $i=1\dots l$, with $x=x_{0}$, we obtain:
@@
Letting $y_{i-1}=D_{i}x_{i-1}+U_{i}y_{i}$, $i=1,\dots,l$, $y_{l}=S_{l}x_{l}$, with $D_{1}=D$ and $y_{0}=y$, we obtain:
@@
For the $\mathcal{H}^{2}$ matrix $A\in\mathbb{R}^{N\times N}$, according to [6], the storage requirement scales as $\mathcal{O}(N)$. $\mathcal{H}^{2}$ construction and matrix-vector multiplication complexity is also $\mathcal{O}(N)$.
@@ 3.2 Multigrid method
A multigrid method [9, 20] is a powerful numerical technique used for solving partial differential equations (PDEs) and linear systems of equations with sparse matrices. While a solver in its own right, it is often used to precondition other iterative algorithms to help them converge rapidly.
@@ 3.3 $\mathcal{H}^{2}$-MG
In this section, we present a multigrid algorithm for a system with an $\mathcal{H}^{2}$ matrix. Consider
@@
In our numerical experiments, which are on symmetric positive definite matrices, we use a CG [36, 25] solver as a smoother because of its fast and easily parallelized application. Thanks to the $\mathcal{H}^{2}$ structure of matrix $A$, the application of one iteration of CG is linear in time and memory cost. After smoothing, we receive an approximation of the error:
@@
We then compute the residual:
$\hat{r}_{1}=r_{1}-A\tilde{e}_{1}$
and subtract the term $A\tilde{e}_{1}$ from both parts of equation (6) to obtain
@@
Letting $\hat{e}_{1}=e_{1}-\tilde{e}_{1}$, we obtain the system
@@
We then build the restriction and prolongation operators. If we write the matrix $A$, with orthogonal bases $U$ and $V$, explicitly in its $\mathcal{H}^{2}$ format, we obtain an expanded version of Equation (7):
@@
Using the orthogonality of $U_{1}$ and $V_{1}$, this may be written as:
@@
We define the restricted operator $A_{2}$ as:
(8) $A_{2}=U^{\top}_{1}DV^{\top}_{1}+D_{2}+U_{2}(\dots(D_{l}+U_{l}S_{l}V_{l})\dots)V_{2}.$
@@
Note that $A_{2}\in\mathbb{R}^{N_{2}\times N_{2}}$ has $\mathcal{H}^{2}$ structure, just like the matrix $A$. The matrices $A$ and $A_{2}$ have exactly the same structure, except that the matrix $A$ stores the full matrix $D$, while $A_{2}$ stores a reduced part of it, $U_{1}^{\top}DV^{\top}$. Thus, matrix $A_{2}$ stores less information than $A$. Since $A$ scales linearly as an $\mathcal{H}^{2}$ matrix, $A_{2}$ also scales linearly. See the detailed proof in Section 3.4.
We also defined the restricted error and residual vectors as:
@@
$A_{2}e_{2}=r_{2}.$
Then, according to the multigrid algorithm, we apply a smoother to the restricted system and obtain an approximation of the error:
@@
where $A_{l}$ is a small dense matrix since $l$ is the top level of $\mathcal{H}^{2}$ hierarchy. We solve the system directly. In our computations, we use the Cholesky factorization:
@@
Then, we apply the smoothing operator starting with the initial guess $\tilde{e}_{l-1}$:
$e_{l-1}=\textbf{Iter}(A_{l-1},r_{l-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tilde{e}_{l-1}}).$
@@
$\tilde{e}_{1}=\textbf{Iter}(A,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}r_{1}},0)$$e_{1}=\textbf{Iter}(A,r_{1},\tilde{e}_{1})$$r_{2}=U_{1}^{\top}(r_{1}-A\tilde{e}_{1})$$r_{3}=U_{2}^{\top}(r_{2}-A_{2}\tilde{e}_{2})$$\tilde{e}_{1}=\tilde{e}_{1}+V_{1}^{\top}e_{2}$$\tilde{e}_{2}=\tilde{e}_{2}+V_{2}^{\top}e_{3}$$\tilde{e}_{2}=\textbf{Iter}(A_{2},r_{2},0)$$e_{2}=\textbf{Iter}(A_{2},r_{2},\tilde{e}_{2})$$e_{3}=\textbf{dir\_sol}(A_{3},r_{3})$
@@
The visualization of $\mathcal{H}^{2}$-MG V-cycle, presented in Figure 4 emphasizes the analogy of the $\mathcal{H}^{2}$-MG method with the $\mathcal{H}^{2}$ structure. Compare $\mathcal{H}^{2}$-MG V-cycle in Figure 4 and the $\mathcal{H}^{2}$ matrix-vector product in Figure 3.